# Predicting Time Pressure of Powered Two-Wheeler Riders for Proactive Safety Interventions

### v1

@@ II-G Research gap and motivation
To address this, our research is structured into four sequential phases. First, we collected a first-of-its-kind, high-resolution PTW simulator dataset comprising more than 129,000 labeled multivariate time-series sequences from 153 rides by 51 participants under three TP conditions (NTP, LTP, HTP). Second, we conduct an empirical analysis of TP Impact to establish how it manifests in rider behaviour, confirming its association with a degradation of steering stability, inadequate compensation for risk, and a systemic escalation of overall risk. Third, we use these behavioural insights to develop a predictive model for TP, directly estimating the cognitive state of sensor-derived signals. Finally, this model establishes a pathway to proactive safety systems under the Safe System Approach, enabling real-time detection of high-risk states to inform future PTW rider-assistance systems, ITS, and safer PTW mobility.
@@ III Methodology for Data Collection
This section outlines the methodology for data collection, which involved a simulator study with specifically designed TP scenarios to observe transitions in rider behavior.
III-A Simulator Setup
All experimental trials are conducted using a static PTW simulator (Technotrove Simulation Pvt. Ltd. [Technotrove]) housed at the Indian Institute of Technology Indore. The simulator is built upon an instrumented Honda motorcycle frame equipped with fully functional throttle, brake, clutch, and gear controls, thereby replicating the ergonomics of actual on-road riding. A three-screen immersive display system surrounds the rider to provide a wide field of view and realistic roadway visualization. The simulator is integrated with a real-time data acquisition system that continuously records rider inputs, vehicle dynamics, and environmental parameters using high-precision industrial-grade sensors with documented accuracy specifications. Complete technical specifications, including sensor accuracy ratings per IEC 393 standards, are provided in Table I. This configuration provides a high-fidelity yet safe experimental environment, allowing systematic investigation of rider responses under various traffic and roadway scenarios.
@@ III-A Simulator Setup
III-B Simulator Scenario Design
@@
III-C Participants
The vast majority of two-wheeler riders in India are male, as reflected in MoRTH road fatality statistics [morth2019, morth2020, morth2021, morth2022, morth2023]. This provides a clear rationale for focusing on male participants in this study. Accordingly, fifty-one male riders (ages 18–42) with at least two years of riding experience participated in the study.
According to the official Road Accidents in India reports by MoRTH, road fatalities are predominantly male. In 2019, males accounted for approximately 86% of all fatalities [morth2019], increasing slightly to 87.3% in 2020 [morth2020]. This pattern persisted in 2021 (86.4% male) [morth2021], 2022 (86.2% male) [morth2022], and 2023 (85.2% male) [morth2023], with the majority of fatalities concentrated in the 18–45 age group across all years. From this, we emphasize that the overwhelming male predominance in two-wheeler fatalities in India provides a strong empirical basis for the male-focused composition of our participant sample. The implications of this sampling strategy for model generalizability are acknowledged and further addressed in Section VIII-A. Detailed rider characteristics are presented in Table II, and Fig. 3 shows participants performing simulator tasks.
@@ III-D TP Scenarios
III-E Behavioral Transition Under Exam Pressure: Why LTP Matters
@@
III-F Dataset Collection and Experimental Design
@@ IV Methodology for TP prediction
This section presents the architecture and mathematical formulation of MTPS
@@ V-B Baseline Models and Evaluation Metrics
MTPS is benchmarked against various ML and DL models: decision tree (DT), support vector machine (SVM), random forest (RF), recurrent neural network (RNN), convolutional neural network (CNN), simplified hyperGraph network (SHGN) [tang2024simplifying], and time series transformer (TST) [10372632]. Performance metrics include classification (Accuracy, F1, Precision, Recall, ROC-AUC) and regression (MAE, MSE, $R^{2}$).
@@ V-C Implementation and Training Details
All models are implemented in Python using PyTorch and Scikit-learn and trained on an NVIDIA T400 GPU (4 GB VRAM). The features are standardized, and data are split into training and testing sets with an 80:20 stratified strategy. The Adam optimizer [kingma2014adam] with sparse categorical cross-entropy loss is used, and hyperparameters are tuned on the validation set. See Table IV for details.
@@
Data Preprocessing Train-Test Split / Stratified / Standardization 80:20 / Yes Zero-Mean, Unit Var.
@@ VI Results and Analysis
This section presents empirical analysis of TP and comprehensive evaluation of the proposed MTPS model against eight state-of-the-art benchmarks to predict rider TP.
@@ VI-A Empirical Analysis of TP Impact on Speed and Control
Our statistical analysis reveals systematic behavioral degradation under TP across multiple riding domains. As quantified in Table V, HTP conditions induced 48% higher mean speeds and 36% greater speed variability compared to NTP. All speed values are in km/h. Brake forces are recorded as normalized outputs (0–100%) from the simulator’s braking sensors, where the 604% increase in front brake application under HTP indicates substantially more aggressive braking behavior. The progressive increase in both mean speed and standard deviation across NTP $\rightarrow$ LTP $\rightarrow$ HTP conditions demonstrates that TP systematically degrades riding control and increases risk exposure. Control patterns showed equally dramatic degradation, with 604% higher front brake force and 35% increased gear usage under HTP conditions.
@@ VI-B The Inadequacy of Rider Compensation Under TP
TP significantly compromises riding stability. As Table VI shows, riders under HTP have 21.3% worse control, yet commit 10.2% more dangerous mistakes compared to the NTP condition. This indicates that even when riders attempt to compensate for degraded control, they cannot fully prevent unsafe events. In this analysis, control variability is measured as the standard deviation of the steering wheel angle (in degrees), and safety violations represent the total count of critical errors such as dangerous overtaking, signal violations, improper gap maintenance and lane-crossing errors—per ride. Values are presented as Mean ± Standard Deviation. These metrics directly quantify how degraded control translates into behaviours that increase crash risk.
@@ VI-C Systemic Risk Escalation Under TP
Consistent with the literature and confirmed by our empirical analysis, TP significantly increases collision risk and control instability, as evidenced by the consistent deviations in speed control, brake application, and clutch management. These findings provide a clear empirical basis for the MTPS framework, demonstrating the need to predict cognitive stress (TP) to enable proactive intervention
@@ VI-D1 Overall Performance Comparison
TABLE VIII: Results comparison of all proposed models including MTPS.
@@
DT 55.03 55.19 56.48 55.03 69.61 0.5553 0.7663 -0.1091
SVM 61.53 61.37 61.36 61.53 78.14 0.5012 0.7343 -0.0627
RF 67.03 67.77 71.82 67.03 86.73 0.3672 0.4421 0.3602
CNN 80.75 80.95 87.25 80.75 95.74 0.1969 0.2057 0.7023
RNN 81.97 82.05 84.25 81.97 95.37 0.2194 0.2975 0.5695
SHGN 84.56 84.55 84.60 84.56 96.51 0.2039 0.1130 0.0000
TST 89.40 89.45 90.42 89.40 98.35 0.1138 0.1294 0.8127
Informer 90.11 90.11 90.10 90.11 98.06 0.1105 0.1336 0.8066
MTPS (Proposed) 91.53 91.54 91.56 91.53 98.93 0.0928 0.1090 0.8423
The proposed MTPS model is benchmarked against eight diverse baseline models spanning traditional machine learning (DT, SVM, RF), deep learning (CNN, RNN), and state-of-the-art architectures (SHGN, TST, Informer). As shown in Table VIII, MTPS achieves SOTA performance across all metrics with 91.53% accuracy and 98.93% ROC AUC, alongside the lowest MAE (0.0928) and highest $R^{2}$ (0.8423). This consistent improved performance across fundamentally different architectural paradigms demonstrates the architectural robustness of MTPS beyond any single modeling approach.
@@ VI-D2 Ablation Study
To validate the architectural design of MTPS, we conducted a comprehensive ablation study by systematically removing key components. The results presented in Table IX demonstrate that each element contributes significantly, with the full integrated architecture achieving optimal results.
@@ VI-D3 Statistical Significance of MTPS
To assess significance, Paired t-tests [HEDBERG2015277] and one-way ANOVA [Zhang2012TrafficSafety] are conducted on five-fold precision scores ($\alpha=0.05$). Results in Table X confirm that MTPS delivers statistically significant gains over all baselines.
@@
DT 729.64 0.000000 Significant
SVM 462.66 0.000000 Significant
RF 590.43 0.000000 Significant
CNN 266.18 0.000000 Significant
RNN 211.45 0.000000 Significant
SHGN 133.21 0.000000 Significant
TST 24.90 0.000015 Significant
Informer 18.40 0.000051 Significant
One-way ANOVA 81704.49 0.000000 Significant
VI-D4 Importance of TP Prediction
The utility of MTPS-predicted TP is demonstrated by its integration into a downstream Informer crash risk model, detailed in Section VII-A. Using predicted TP as a feature increases collision prediction accuracy from 91.25% to 93.51%, approaching the performance of an oracle model with ground-truth TP (93.72%). This shows that predicted TP, which captures the cognitive why behind risky maneuvers, provides a unique signal for proactive safety systems that kinematics alone cannot offer, underscoring its critical role in rider safety management.
VI-D5 Impact of Feature Subsets on MTPS Performance
@@ VI-E5 Model Performance and Confidence Intervals
Class-wise precision, recall, and F1 score are computed directly from the confusion matrix (Fig. 6). 95% confidence intervals (CIs) are computed using the Wilson score interval [rajaraman2022deep, binomialciwilsonwikipedia]:
$CI_{95\%}=\frac{p+\frac{z^{2}}{2n}\pm z\sqrt{\frac{p(1-p)}{n}+\frac{z^{2}}{4n^{2}}}}{1+\frac{z^{2}}{n}},\quad z=1.96,$ (12)
where $p$ is the observed proportion (precision or recall) and $n$ is the sample size and $z=1.96$ for 95% confidence  [binomialciwilsonwikipedia].
@@
VI-E6 Worked Example: Calm / NTP Class
This section shows how we obtain calculations that directly support the Calm (NTP) threshold (Table XII and Fig. 8) assignment ($P(\text{NTP})\geq 0.80$, $P(\text{LTP})\leq 0.15$, $P(\text{HTP})\leq 0.10$). From the confusion matrix (Fig. 6), we computed the values: $TP=7277$, $FP=140+566=706$, $FN=34+1291=1325$. $\text{Precision}=\frac{7277}{7277+706}=0.912,\;\;\text{Recall}=\frac{7277}{7277+1325}=0.846$, $F1\approx 0.878$. Applying Eq.(12) where $n$ is the total relevant sample for the metric: for precision, $n=TP+FP=7277+706=7983$; for recall, $n=TP+FN=7277+1325=8602$. We obtain $\text{Precision CI}\approx[0.905,0.918],\;\;\text{Recall CI}\approx[0.838,0.853]$.
The class-wise performance of the MTPS with the Wilson CIs 95% is shown in Table XIII. This ensures stable baseline detection while minimizing false alarms, providing a statistically justified basis for Table XII.
@@ VI-F1 Threshold Selection and Statistical Support
The threshold values in Table XII were empirically determined through a multi-faceted analysis. First, we identified calibration-consistent regions in Fig. 7b where predicted probabilities align with observed frequencies. Second, we analyzed precision-recall trade-offs using confusion matrix metrics to balance safety-critical detection rates with false alarm minimization. Finally, we incorporated human observation to ensure practical relevance for rider assistance systems. This integrated approach ensures the thresholds are statistically supported and reliable for both baseline monitoring and safety-critical ITS decision-making. Grounded in threshold calibration theory [rajaraman2022deep, 10.5555/3618408.3619385], this framework leverages calibrated probabilities for robust risk stratification [9137689, 1637672]. For PTWs, where perception and urgency strongly shape warning effectiveness, these dynamically calibrated thresholds enhance both safety and interpretability [Carrington_2023].
@@ VI-G MTPS Model Complexity and Edge-AI Deployability
TABLE XIV: Metrics for the MTPS model predicting rider TP.
@@
Model Size 2.16 MB Lightweight; easy to store and deploy on most devices
@@
Target Deployment Regions Global Can operate on wearables, on- board units and in low-resource settings
@@ VII-A1 Evaluation of MTPS-Predicted TP
The Informer model [zhou2021informer], a Transformer-based architecture for long-sequence time-series forecasting, is used to assess the utility of MTPS-predicted TP in collision prediction. Given a multivariate input $X=\{x_{1},\ldots,x_{T}\},x_{t}\in\mathbb{R}^{d}$, it learns
@@
TABLE XV: Informer model performance with different TP inputs for collision prediction
Input Configuration Accuracy (%)
@@
As shown in Table XV, incorporating MTPS-predicted TP into an Informer-based crash risk predictor increases accuracy from 91.25% (baseline) to 93.51%, approaching the oracle performance of 93.72%. The Oracle establishes the theoretical upper bound with $\Delta_{\text{max}}=\mathcal{A}_{\text{oracle}}-\mathcal{A}_{\text{baseline}}=2.47\%$. MTPS-predicted TP achieves $\mathcal{A}_{\text{MTPS}}=93.51\%$, yielding $\Delta_{\text{MTPS}}=\mathcal{A}_{\text{MTPS}}-\mathcal{A}_{\text{baseline}}=2.26\%$ and a minimal gap $\epsilon=|\mathcal{A}_{\text{oracle}}-\mathcal{A}_{\text{MTPS}}|=0.21\%$ to the Oracle ($\mathcal{A}_{\text{oracle}}=93.72\%$), corresponding to $\eta=\tfrac{\Delta_{\text{MTPS}}}{\Delta_{\text{max}}}\times 100\%\approx 91.50\%$ of the maximum benefit. TP can be interpreted as a latent variable influencing rider dynamics. The inference pipeline $\mathbf{X}\xrightarrow{f_{\text{MTPS}}}\hat{y}_{\text{MTPS}}\xrightarrow{\oplus}X\oplus\hat{y}_{\text{MTPS}}\xrightarrow{f_{\text{Informer}}}\hat{y}$ shows that predicted TP captures over $91.50\%$ of the ground-truth benefit. This demonstrates that TP effectively reflects rider cognitive stress and control variability, serving as a discriminative, interpretable feature for real-time ITS interventions.

### v3

@@ II-G Research Gap and Motivation
To address this, our research is structured into four sequential phases. First, we collected a first-of-its-kind, high-resolution PTW simulator dataset comprising more than 129,000 labeled sequences from 153 rides by 51 participants under three TP conditions (NTP, LTP, HTP). Second, we conducted an empirical analysis of TP impact to establish how it manifests in rider behavior. Third, we developed MTPS to predict TP from sensor data. Finally, we established a pathway to proactive safety systems under the Safe System Approach (SSA) using MTPS-predicted TP for collision prediction and threshold-based rider states for graded ITS interventions.
@@ III Methodology for Data Collection
Based on the gaps identified in the literature review, we designed a controlled simulator study to collect high-resolution PTW riding data under varying TP conditions. This section describes the experimental setup, participant recruitment, TP scenarios, and data collection protocol.
III-A Experimental Design
III-A1 Simulator Setup
All experimental trials are conducted using a static PTW simulator (Technotrove Simulation Pvt. Ltd. [33]) housed at the Indian Institute of Technology Indore. The simulator is built upon an instrumented Honda motorcycle frame equipped with fully functional throttle, brake, clutch, and gear controls, thereby replicating the ergonomics of actual on-road riding. A three-screen immersive display system surrounds the rider to provide a wide field of view and realistic roadway visualization. The simulator is integrated with a real-time data acquisition system that continuously records rider inputs, vehicle dynamics, and environmental parameters using high-precision industrial-grade sensors with documented accuracy specifications. Complete technical specifications, including sensor accuracy ratings per IEC 393 standards, are provided in Table I. This configuration provides a high-fidelity yet safe experimental environment, allowing systematic investigation of rider responses under various traffic and roadway scenarios.
@@ III-A1 Simulator Setup
III-A2 Simulator Scenario Design
@@
III-A3 Participants
The vast majority of two-wheeler riders in India are male, as reflected in MoRTH road fatality statistics [21, 22, 23, 24, 25]. According to official MoRTH reports, males accounted for 86.0%, 87.3%, 86.4%, 86.2%, and 85.2% of PTW fatalities between 2019 and 2023, respectively, with most deaths in the 18–45 age group. This overwhelming male predominance provides a strong empirical basis for focusing on male participants in this study. Accordingly, fifty-one male riders (ages 18–42) with at least two years of riding experience participated. Detailed rider characteristics are presented in Table II, and Fig. 3 shows participants performing simulator tasks.
@@ III-A4 TP Scenarios
III-A5 Behavioral Transition Under Exam Pressure: Why LTP Matters
@@
III-B Data Collection
III-B1 Dataset Collection and Experimental Design
@@ III-B2 Dataset Details
Collecting real-world TP data is challenging due to safety risks and ethical constraints, but we will work within these constraints. We plan to collaborate with commercial riding platforms to collect naturalistic riding data and validate MTPS in real traffic conditions as part of our future work.
@@ IV Methodology for TP prediction
This section presents the architecture and mathematical formulation of MTPS.
@@ V-B Baseline Models and Evaluation Metrics
MTPS is benchmarked against various ML and DL models: decision tree (DT), support vector machine (SVM), random forest (RF), recurrent neural network (RNN), convolutional neural network (CNN), simplified hyperGraph network (SHGN) [42], and time series transformer (TST) [43]. Performance metrics include classification (Accuracy, F1, Precision, Recall, ROC-AUC) and regression (MAE, MSE, $R^{2}$).
@@ V-C Implementation and Training Details
All models are implemented in Python using PyTorch and Scikit-learn and trained on an NVIDIA T400 GPU (4 GB VRAM). The features are standardized, and data are split into training, testing and validation sets with an 80:10:10 stratified strategy. The Adam optimizer [14] with sparse categorical cross-entropy loss is used, and hyperparameters are tuned on the validation set. See Table IV for details.
@@
Data Preprocessing Train-Test-Val Split / Stratified / Standardization 80/10/10 / Yes / Zero-Mean, Unit Var.
@@ VI Results and Analysis
We present our results in three parts. First, we analyze how TP systematically degrades riding behavior. Second, we evaluate MTPS’s classification performance against six baselines. Third, we demonstrate the practical utility of predicted TP for collision prediction and threshold-based ITS interventions.
@@ VI-A Empirical Analysis of TP Impact on Speed and Control
Our statistical analysis reveals systematic behavioral degradation under TP across multiple riding domains, consistent with earlier findings by Pawar et al. [32] for four-wheelers, but extended here to PTW riders. As quantified in Table V, HTP conditions induced 48% higher mean speeds and 36% greater speed variability compared to NTP. All speed values are in km/h. Brake forces are recorded as normalized outputs (0–100%) from the simulator’s braking sensors, where the 604% increase in front brake application under HTP indicates substantially more aggressive braking behavior. The progressive increase in both mean speed and standard deviation across NTP $\rightarrow$ LTP $\rightarrow$ HTP conditions demonstrates that TP systematically degrades riding control and increases risk exposure. Control patterns showed equally dramatic degradation, with 604% higher front brake force and 35% increased gear usage under HTP conditions.
@@ VI-B The Inadequacy of Rider Compensation Under TP
TP significantly compromises riding stability. As shown in Table VI, riders under HTP have 21.3% worse control, yet commit 10.2% more dangerous mistakes compared to the NTP condition. This indicates that even when riders attempt to compensate for degraded control, they cannot fully prevent unsafe events. In this analysis, control variability is measured as the standard deviation of the steering wheel angle (in degrees), and safety violations represent the total count of critical errors such as dangerous overtaking, signal violations, improper gap maintenance and lane-crossing errors—per ride. Values are presented as Mean ± Standard Deviation. These metrics directly quantify how degraded control translates into behaviours that increase crash risk.
@@ VI-C Systemic Risk Escalation Under TP
Consistent with the literature and confirmed by our empirical analysis, TP significantly increases collision risk and control instability, as evidenced by the consistent deviations in speed control, brake application, and clutch management. These findings provide a clear empirical basis for the MTPS framework, demonstrating the need to predict cognitive stress (TP) to enable proactive intervention.
@@ VI-D1 Overall Performance Comparison
TABLE VIII: Results comparison of all proposed models including MTPS. We evaluate each model five times and report the mean (standard deviation).
@@
DT 79.21 ± 0.58 79.50 81.68 79.21 93.83 0.2291 0.2716 0.6070
RNN 81.97 ± 0.49 82.05 84.25 81.97 95.37 0.2194 0.2975 0.5695
CNN 85.61 ± 0.54 85.52 85.95 85.61 97.21 0.1678 0.2155 0.6881
SHGN 89.33 ± 0.45 89.41 90.26 89.33 98.27 0.1107 0.1188 0.8281
TST 89.40 ± 0.28 89.45 90.42 89.40 98.35 0.1138 0.1294 0.8127
Informer 90.11 ± 0.39 90.11 90.10 90.11 98.06 0.1105 0.1336 0.8066
MTPS (Proposed) 91.53 ± 0.27 91.54 91.56 91.53 98.93 0.0928 0.1090 0.8423
The proposed MTPS model is benchmarked against six diverse baseline models spanning traditional machine learning, deep learning, and state-of-the-art architectures (SHGN, TST, Informer). As shown in Table VIII, MTPS achieves SOTA performance across all metrics with 91.53% accuracy and 98.93% ROC AUC, alongside the lowest MAE (0.0928) and highest $R^{2}$ (0.8423). This consistent improved performance across fundamentally different architectural paradigms demonstrates the architectural robustness of MTPS beyond any single modeling approach.
@@ VI-D2 Ablation Study
To validate the architectural design of MTPS, we conducted a comprehensive ablation study by systematically removing key components. The results presented in Table IX demonstrate that each element contributes significantly, with the full integrated architecture achieving optimal results. Removing attention causes the largest drop (0.81%), followed by SE blocks (0.64%), while Conv-only reduces accuracy by 1.11%.
@@ VI-D3 Statistical Significance of MTPS
To assess significance, paired t-tests [12] and one-way ANOVA [47] are conducted on five-fold accuracy scores ($\alpha=0.05$). Results in Table X confirm that MTPS delivers statistically significant gains over all baselines. Paired t-tests between MTPS and each baseline (DT, RNN, CNN, SHGN, TST, Informer) yield p-values <0.001<0.001 and t-statistics from $14.47$ (vs. Informer) to $61.60$ (vs. DT). The one-way ANOVA comparing all models collectively produces an F-statistic of $3812.78$ with $p<0.000001$, confirming that the observed differences in model performance are statistically significant overall. Additionally, Cohen’s d effect sizes range from $9.52$ to $73.62$, indicating very large practical significance. These results robustly confirm the statistical significance of MTPS for time pressure classification.
@@
DT 61.60 0.000000 Yes
RNN 58.39 0.000000 Yes
CNN 55.72 0.000000 Significant
SHGN 31.76 0.000006 Yes
TST 31.27 0.000006 Yes
Informer 14.47 0.000133 Yes
One-way ANOVA 3812.78 0.000000 Yes
VI-D4 Impact of Feature Subsets on MTPS Performance
@@ VI-E5 Model Performance and Confidence Intervals
Class-wise precision, recall, and F1 score are computed directly from the confusion matrix (Fig. 6). 95% confidence intervals (CIs) are computed using the Wilson score interval [34, 45]:
@@
VI-E6 Worked Example: Calm (NTP) Class
We obtain calculations that support the Calm (NTP) threshold (Table XII and Fig. 8) assignment ($P(\text{NTP})\geq 0.80$, $P(\text{LTP})\leq 0.15$, $P(\text{HTP})\leq 0.10$). The precision, recall and confidence intervals are derived from the confusion matrix (Fig. 6) using 95% CI. The confusion matrix yields $TP=7277$, $FP=706$, $FN=1325$, giving precision =0.912=0.912, recall =0.846=0.846, and F1 =0.878=0.878. The class-wise performance of the MTPS with the Wilson CIs 95% is shown in Table XIII. This ensures stable baseline detection while minimizing false alarms, providing a statistically justified basis for Table XII.
@@ VI-F1 Threshold Selection and Statistical Support
The threshold values in Table XII were empirically determined through a multi-faceted analysis. First, we identified calibration-consistent regions in Fig. 7b where predicted probabilities align with observed frequencies. Second, we analyzed precision-recall trade-offs using confusion matrix metrics to balance safety-critical detection rates with false alarm minimization. Finally, we incorporated human observation to ensure practical relevance for rider assistance systems. This integrated approach ensures the thresholds are statistically supported and reliable for both baseline monitoring and safety-critical ITS decision-making. Grounded in threshold calibration theory [34, 15], this framework leverages calibrated probabilities for robust risk stratification. For PTWs, where perception and urgency strongly shape warning effectiveness, these dynamically calibrated thresholds enhance both safety and interpretability [13].
@@ VI-G MTPS Model Complexity and Edge-AI Deployability
TABLE XIV: Model complexity and Edge-AI deployability of MTPS.
@@
Inference Time 0.21 ms (i7 CPU) Supports sub-100ms real-time intervention on edge devices
Model Size 0.66 MB Lightweight; easy to store and deploy on most devices
@@
Deployment Regions Global Can operate on wearables, on- board units and in low-resource settings
@@ VII-A1 Evaluation of MTPS-Predicted TP
The Informer model [49], a Transformer-based architecture for long-sequence time-series forecasting, is used to assess the utility of MTPS-predicted TP in collision prediction. Given a multivariate input $X=\{x_{1},\ldots,x_{T}\},x_{t}\in\mathbb{R}^{d}$, it learns
@@
TABLE XV: Collision prediction accuracy of Informer and TimesNet with baseline, MTPS-predicted TP, and oracle.
Model and Input Configuration Accuracy (%)
Informer
@@
TimesNet
Baseline: $f_{\text{TimesNet}}(X)$ 92.10
MTPS-Predicted TP: $f_{\text{TimesNet}}(X\oplus\hat{y}_{\text{MTPS}})$ 93.90
Oracle (GT TP): $f_{\text{TimesNet}}(X\oplus y)$ 94.06
As shown in Table XV, incorporating MTPS-predicted TP into an Informer-based crash risk predictor increases accuracy from 91.25% (baseline) to 93.51%, approaching the oracle performance of 93.72%. The Oracle establishes the theoretical upper bound with $\Delta_{\text{max}}=\mathcal{A}_{\text{oracle}}-\mathcal{A}_{\text{baseline}}=2.47\%$. MTPS-predicted TP achieves $\mathcal{A}_{\text{MTPS}}=93.51\%$, yielding $\Delta_{\text{MTPS}}=\mathcal{A}_{\text{MTPS}}-\mathcal{A}_{\text{baseline}}=2.26\%$ and a minimal gap $\epsilon=|\mathcal{A}_{\text{oracle}}-\mathcal{A}_{\text{MTPS}}|=0.21\%$ to the Oracle ($\mathcal{A}_{\text{oracle}}=93.72\%$), corresponding to $\eta=\tfrac{\Delta_{\text{MTPS}}}{\Delta_{\text{max}}}\times 100\%\approx 91.50\%$ of the maximum benefit. TP can be interpreted as a latent variable influencing rider dynamics. The inference pipeline $\mathbf{X}\xrightarrow{f_{\text{MTPS}}}\hat{y}_{\text{MTPS}}\xrightarrow{\oplus}X\oplus\hat{y}_{\text{MTPS}}\xrightarrow{f_{\text{Informer}}}\hat{y}$ shows that predicted TP captures over $91.50\%$ of the ground-truth benefit. The generalizability of this finding is further supported by TimesNet [46], which achieved 94.06% accuracy with ground truth TP (93.90% with MTPS-predicted TP), compared to 92.10% without TP. This demonstrates that TP effectively reflects rider cognitive stress and control variability, serving as a discriminative, interpretable feature for real-time ITS interventions.